The firm receives an average of $20,000 in checks per day. The weighted average delay in clearing the checks received is 3 days. Meanwhile, the firm writes an average of $17,000 in checks to pay its suppliers per day. The usual clearing time for the checks the firm wrote is 2 days. The current interest rate is 0.015 percent per day. What is the most the firm should be willing to pay today (in a lump sum today) to eliminate its float entirely?

A) 3000

B) 26000

C) 34000

D) 37000

E) 60000
Answer #1

Firm's Collection Float = Delay in Check Clearing * Average Amount = 3 * 20,000 = 60,000

Firm's Disbursement Float = Delay in Check Clearing * Average Amount = 2 * 17,000 = 34,000

Net float = 60000-34000=26000

So, option B is correct.